find the general form of the equation of the line that passes through the points (-7,-1) and (6,7)
Answer by mananth(16946)

x1 y1 x2 y2
-7 -1 6 7

slope m = (y2-y1)/(x2-x1)
( 7 - -1 )/( 6 - -7 )
( 8 / 13 )
m= 8/13

Plug value of the slope and point ( -7,-1) in
Y = m x + b
-1.00 = -63/7 + b
b= -1 - 30/7
b= -37/13
So the equation will be
Y = 5/8 x -37/7
LCD = 56
multiply by 56
56y=35x-296
35x-56y=296
m.ananth@hotmail.ca